# reference.bib

### v1

@@ 3. Fourier-Jacobi Forms
In this section, we will define Fourier-Jacobi forms of lattice index. We mainly follow the thesis in [jacobi_lattice] and Krieg in [krieg_jacobi].
@@
The group $\hbox{SL}_{2}(\mathbb{R})$ acts on $H^{(\Lambda,\sigma)}(\mathbb{R})$ from the right, via
@@
The real Jacobi group associated with $(\Lambda,\sigma)$, denoted with $J^{(\Lambda,\sigma)}(\mathbb{R})$, is defined to be the semi-direct product of $\hbox{SL}_{2}(\mathbb{R})$ and $H^{(\Lambda,\sigma)}(\mathbb{R})$. The composition law is then
@@
We also define the integral Jacobi group to be the semi-direct product of $\hbox{SL}_{2}(\mathbb{Z})$ and $H^{(\Lambda,\sigma)}(\mathbb{Z})$ and we will denote it by $J^{(\Lambda,\sigma)}$.
We are now going to define some slash operators, acting on holomorphic, complex-valued functions on ×$\mathbb{H}\times(\Lambda\otimes\mathbb{C})$.
@@
Let $k$ be a positive integer and $f:\mathbb{H}\times(\Lambda\otimes\mathbb{C})\longrightarrow\mathbb{C}$ a holomorphic function. For $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\hbox{SL}_{2}(\mathbb{R})$ , we define:
@@
In the case when $M\in\hbox{GL}_{2}^{+}(\mathbb{R})$, we use $\det(M)^{-1/2}M$ instead of $M$. For $h=(x,y,\zeta)\in H^{(\Lambda,\sigma)}(\mathbb{R})$:
@@
Finally, for the action of $J^{(\Lambda,\sigma)}(\mathbb{R})$ on complex-valued, holomorphic functions on ×$\mathbb{H}\times(L\otimes\mathbb{C})$, we have:
@@
  • For all $\gamma\in J^{(\Lambda,\sigma)}$, we have
@@
    $f(\tau,z)=\sum_{n\in\mathbb{Z},r\in\Lambda^{*},2n\geq\sigma[r]}c_{f}(n,r)e(n\tau+\sigma(r,z)).$
We denote the space of such forms by $J_{k}(\Lambda,\sigma)$. We say $f$ is a Jacobi cusp form if $c_{f}(n,r)=0$, when $2n=\sigma[r]$. We denote the space of Fourier-Jacobi cusp forms by $S_{k}(\Lambda,\sigma)$.
@@
Let $\phi,\psi\in S_{k}(\Lambda,\sigma)$. If then $U\leq J^{(\Lambda,\sigma)}$ of finite index, we define their Petersson inner product via:
$\langle\phi,\psi\rangle_{U}=\frac{1}{\left[J^{\Lambda}:U\right]}\int_{U\backslash\mathcal{H}\times(\Lambda\otimes\mathbb{C})}\phi(\tau,z)\overline{\psi(\tau,z)}v^{k}e^{-2\pi\sigma(y,y)v^{-1}}\hbox{d}V,$
where $\tau=u+iv$, $z=x+iy$ and $\hbox{d}V:=v^{-n-2}\hbox{d}u\hbox{d}v\hbox{d}x\hbox{d}y$. This inner product does not depend on the choice of $U$, so in what follows, we drop the subscript.
We now specify to our case by taking $\Lambda=L=\mathbb{Z}^{n}$ and $\sigma(x,y)=x^{t}Sy$ for all $x,y\in V$. We also write $J_{S}$ for the intergal Jacobi group in this case, i.e. $J_{S}=\hbox{SL}_{2}(\mathbb{Z})\rtimes(\mathbb{Z}^{n}\times\mathbb{Z}^{n})$.
@@
$\Gamma_{S,J}:=\left\{\begin{pmatrix}*&*\\ 0&D\end{pmatrix}\in\Gamma_{S}\mid D\in\mathbb{Z}^{2\times 2}\right\}.$
@@
with $D^{*}:=D\left[\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}\right]$ (cf. [eisenstein_thesis, page 44]). We denote by $\Gamma_{S,J}^{\bullet}:=\iota(J_{S})$. Then, by [ajouz, Proposition 2.2.7], we have that the action of $J_{S}$ on ×$\mathbb{H}\times\mathbb{C}^{n}$ is given by
@@
where $j^{\prime}$ denotes the usual factor of automorphy for the action of $\hbox{SL}_{2}(\mathbb{Z})$ on $\mathbb{H}$.
@@
is the same as the action of $J_{S}$ on ×$\mathbb{H}\times\mathbb{C}^{n}$. We now have the following Proposition regarding fundamental domains.
@@
$\mathcal{F}^{J}=\left\{Z=(\omega,z,\tau)\in\mathcal{H}_{S}:(z,\tau)\in\mathcal{F},y_{1}y_{2}-\frac{1}{2}S[v]>0,-\frac{1}{2}\leq x_{1}\leq\frac{1}{2}\right\},$
where $\mathcal{F}$ is a fundamental domain of the action of $J_{S}$ on ×$\mathbb{H}\times\mathbb{C}^{n}$.
@@
Let $Z=(\omega,z,\tau)\in\mathcal{H}_{S}$. We can then pick $g\in\Gamma_{S,J}^{\bullet}$ such that $\iota^{-1}(g)\in J_{S}$ and that $Z^{\prime}:=g\langle Z\rangle=(\omega^{\prime},z,\tau)$ with $(z,\tau)\in\mathcal{F}$ and $\omega^{\prime}\in\mathbb{H}$ arbitrary. This follows by the fact that the actions are the same, as we have shown above. If now
@@
for $\lambda\in\mathbb{Z}^{n+2}$, then $T_{\lambda}\in\Gamma_{S}$ (cf. [eisenstein_thesis, Theorem 1.4.4]). We then have $T_{\lambda}\langle Z^{\prime}\rangle=Z^{\prime}+\lambda$ and so we can act with a suitable $T_{\lambda}$, so that $-1/2\leq x_{1}\leq 1/2$. Finally, the other condition follows by the way $\mathcal{H}_{S}$ has been defined. ∎
4. Integral Representation
In this section, we define the Dirichlet series of interest and give its integral representation, in terms of an Eisenstein series of Klingen type.
@@ 4. Integral Representation
$\mathcal{D}_{F,G}(s)=\sum_{m=1}^{\infty}\langle\phi_{m},\psi_{m}\rangle m^{-s}.$
@@
$\mathcal{D}_{F,G}(s)$ converges absolutely for $\textup{Re}(s)>k+1$ and represents a holomorphic function in this domain.
@@
with the constant depending only on $F,G$. Indeed, fix $(\tau,z)\in\mathbb{H}\times\mathbb{C}^{n}$ and write $\tau=u+iv,\textup{ }z=x+iy$. If $G(q)=\sum_{N=1}^{\infty}\phi_{N}(\tau,z)q^{N}$, with $q=e^{2\pi i\omega}$, we have by Cauchy’s integral formula that
[…]
@@
[…]
for any $v^{\prime}>Q[y]/v$. But now $\left|F(Z)\right|\left(Q_{0}[\textup{Im}Z]\right)^{k/2}$ is bounded on $\mathcal{H}_{S}$ from [hauffe21, Lemma 3.28], say by a constant $C>0$. Therefore, after choosing $v^{\prime}=Q[y]/v+1/N$, we have
[…]
@@
$E(Z,s)=\sum_{\gamma\in\Gamma_{S,J}\backslash\Gamma_{S}}\left(\frac{Q_{0}[\textup{Im}(\gamma Z)]}{\textup{Im}((\gamma Z)_{2})}\right)^{s},$
where for $Z=(\omega,z,\tau)\in\mathcal{H}_{S}$, we write $Z_{2}=\tau$.
@@
$\gamma=\begin{pmatrix}*&*\\ 0&D\end{pmatrix},\textup{ }D\in\hbox{SL}_{2}(\mathbb{Z})$
(cf. [krieg_integral_orthogonal, Proposition 3]). Now, for $Z=(\omega,z,\tau)\in\mathcal{H}_{S}$, we have $\gamma\langle Z\rangle_{2}=D\langle\tau\rangle$, where the action on the right denotes the usual action of $\hbox{SL}_{2}(\mathbb{Z})$ on $\mathbb{H}$ (cf. [eisenstein_thesis, page 115]). By [Borcherds_Products, Lemma 3.20], we have
@@
where again $j^{\prime}$ denotes the usual factor of automorphy for the action of $\hbox{SL}_{2}(\mathbb{Z})$ on $\mathbb{H}$. But we have $j(\gamma,Z)=j^{\prime}(D,\tau)$ and hence $E(Z,s)$ is well-defined.
@@
But by [eisenstein_thesis, Theorem 3.1.1], the sum
@@
For $Z\in\mathcal{H}_{S}$ and $s\in\mathbb{C}$ with $\textup{Re}(s)>n+2$, we have
$\langle F(Z)E(Z,s),G(Z)\rangle=\frac{1}{\#\hbox{SO}(S;\mathbb{Z})}(4\pi)^{-(s+k-n-1)}\Gamma(s+k-n-1)\mathcal{D}_{F,G}(s+k-n-1),$
@@
$\hbox{SO}(S;\mathbb{Z}):=\{g\in\hbox{SL}_{n}(\mathbb{Z})\mid g^{t}Sg=S\},$
@@
Let $I(s):=\langle F(Z)E(Z,s)\textup{, }G(Z)\rangle$. If we denote by $\mathcal{Q}$ a fundamental domain for the action of $\Gamma_{S}$ on $\mathcal{H_{S}}$, by using the standard unfolding argument, we have for $\hbox{Re}(s)>n+1$
$\displaystyle I(s)$ $\displaystyle=\int_{\mathcal{Q}}F(Z)\overline{G(Z)}\sum_{\gamma\in\Gamma_{S,J}\setminus\Gamma_{S}}\left(\frac{Q_{0}[\textup{Im}(\gamma Z)]}{\textup{Im}((\gamma Z)_{2})}\right)^{s}\left(Q_{0}[\textup{Im}Z]\right)^{k}\hbox{d}^{*}Z$
$\displaystyle=\frac{1}{[\Gamma_{S,J}:\Gamma_{S,J}^{\bullet}]}\int_{\mathcal{F}^{J}}F(Z)\overline{G(Z)}\left(\frac{Q_{0}[Y]}{\textup{Im}(\tau)}\right)^{s}\left(Q_{0}[Y]\right)^{k-n-2}\hbox{d}X\hbox{d}Y,$
where $\mathcal{F}^{J}$ is a fundamental domain for the action of $\Gamma_{S,J}^{\bullet}$ on $\mathcal{H}_{S}$ and $Z=X+iY$. We note here
$[\Gamma_{S,J}:\Gamma_{S,J}^{\bullet}]=\#\hbox{SO}(S;\mathbb{Z})<\infty,$
because $S$ is positive definite (cf. [eisenstein_thesis, Section 1.7]). Hence, from Proposition 3.2, with the notation the same as there, we get
$I(s)=\frac{1}{\#\hbox{SO}(S;\mathbb{Z})}\int_{\mathcal{F}}\int_{y_{1}y_{2}-\frac{1}{2}S[v]>0}\int_{-1/2\leq x_{1}\leq 1/2}F(Z)\overline{G(Z)}y_{2}^{-s}\left(y_{1}y_{2}-\frac{1}{2}S[v]\right)^{k-n-2+s}\hbox{d}X\hbox{d}Y,$
where $\mathcal{F}$ is a fundamental domain of the action of $J_{S}$ on ×$\mathbb{H}\times\mathbb{C}^{n}$.
We now write $F(Z)=\sum_{m=1}^{\infty}\phi_{m}(\tau,z)e^{2\pi im\omega}$ and $G(Z)=\sum_{m=1}^{\infty}\psi_{m}(\tau,z)e^{2\pi im\omega}$. Using the fact that
$\int_{-1/2}^{1/2}e^{2\pi i(m_{1}-m_{2})x_{1}}\hbox{d}x_{1}=\begin{cases}1&\textup{ if }m_{1}=m_{2}\\ 0&\textup{ otherwise}\end{cases},$
@@
$I(s)=\frac{1}{\#\hbox{SO}(S;\mathbb{Z})}\int_{\mathcal{F}}\int_{y_{1}y_{2}-\frac{1}{2}S[v]>0}\sum_{m=1}^{\infty}\phi_{m}(\tau,z)\overline{\psi_{m}(\tau,z)}e^{-4\pi my_{1}}y_{2}^{-s}\left(y_{1}y_{2}-\frac{1}{2}S[v]\right)^{k-n-2+s}\hbox{d}X\hbox{d}Y.$
@@
$I(s)=\frac{1}{\#\hbox{SO}(S;\mathbb{Z})}\int_{\mathcal{F}}\int_{t=0}^{\infty}\sum_{m=1}^{\infty}\phi_{m}(\tau,z)\overline{\psi_{m}(\tau,z)}e^{-4\pi mt}e^{-2\pi m\frac{S[v]}{y_{2}}}y_{2}^{-s}\left(ty_{2}\right)^{k-n-2+s}\hbox{d}t\hbox{d}u\hbox{d}v\hbox{d}x_{2}\hbox{d}y_{2}.$
@@
[…]
Moreover, in this case, the inner product in 3.5 reads as:
$\langle\phi_{m},\psi_{m}\rangle=\int_{\mathcal{F}}\phi_{m}(\tau,z)\overline{\psi_{m}(\tau,z)}y^{k-n-2}e^{-\frac{2\pi m}{y}S[v]}\hbox{d}u\hbox{d}v\hbox{d}x_{2}\hbox{d}y_{2}.$
@@
$I(s)=\frac{1}{\#\hbox{SO}(S;\mathbb{Z})}(4\pi)^{-(s+k-n-1)}\Gamma(s+k-n-1)\mathcal{D}_{F,G}(s+k-n-1),$
@@
$\langle F(Z)E(Z,s),G(Z)\rangle=\frac{1}{\#\hbox{SO}(S;\mathbb{Z})}(4\pi)^{-(s+k-n-1)}\Gamma(s+k-n-1)\mathcal{D}_{F,G}(s+k-n-1),$

### v2

@@ 3. Fourier-Jacobi Forms
In this Section, we will define Fourier-Jacobi forms of lattice index. We mainly follow Mocanu’s thesis in [jacobi_lattice] and Krieg in [krieg_jacobi].
@@
The group $\textup{SL}_{2}(\mathbb{R})$ acts on $H^{(\Lambda,\sigma)}(\mathbb{R})$ from the right, via
@@
The real Jacobi group associated with $(\Lambda,\sigma)$, denoted by $J^{(\Lambda,\sigma)}(\mathbb{R})$, is defined to be the semi-direct product of $\textup{SL}_{2}(\mathbb{R})$ and $H^{(\Lambda,\sigma)}(\mathbb{R})$. The composition law is then
@@
We also define the integral Jacobi group to be the semi-direct product of $\textup{SL}_{2}(\mathbb{Z})$ and $H^{(\Lambda,\sigma)}(\mathbb{Z})$ and we will denote it by $J^{(\Lambda,\sigma)}$.
We are now going to define some slash operators, acting on holomorphic, complex-valued functions on $\mathbb{H}\times(\Lambda\otimes\mathbb{C})$.
@@
Let $k$ be a positive integer and $f:\mathbb{H}\times(\Lambda\otimes\mathbb{C})\longrightarrow\mathbb{C}$ a holomorphic function. For $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\textup{SL}_{2}(\mathbb{R})$ , we define:
@@
In the case when $M\in\textup{GL}_{2}^{+}(\mathbb{R})$, we use $\det(M)^{-1/2}M$ instead of $M$. For $h=(x,y,\zeta)\in H^{(\Lambda,\sigma)}(\mathbb{R})$:
@@
Finally, for the action of $J^{(\Lambda,\sigma)}(\mathbb{R})$ on complex-valued, holomorphic functions on $\mathbb{H}\times(L\otimes\mathbb{C})$, we have:
@@
  • For all $\gamma\in J^{(\Lambda,\sigma)}$, and $(\tau,z)\in\mathbb{H}\times V_{\mathbb{C}}$, we have
@@
    $f(\tau,z)=\sum_{m\in\mathbb{Z},r\in\Lambda^{*},2m\geq\sigma[r]}c_{f}(m,r)e(m\tau+\sigma(r,z)).$
We denote the space of such forms by $J_{k}(\Lambda,\sigma)$. We say $f$ is a Jacobi cusp form if $c_{f}(m,r)=0$, when $2m=\sigma[r]$. We denote the space of Fourier-Jacobi cusp forms by $S_{k}(\Lambda,\sigma)$.
@@
Let $\phi,\psi\in S_{k}(\Lambda,\sigma)$. If $U\leq J^{(\Lambda,\sigma)}$ of finite index, we define their Petersson inner product via:
$\langle\phi,\psi\rangle_{U}:=\frac{1}{\left[J^{\Lambda}:U\right]}\int_{U\backslash\mathcal{H}\times(\Lambda\otimes\mathbb{C})}\phi(\tau,z)\overline{\psi(\tau,z)}v^{k}e^{-2\pi\sigma(y,y)v^{-1}}\textup{d}V,$
where $\tau=u+iv$, $z=x+iy$ and $\textup{d}V:=v^{-n-2}\textup{d}u\textup{d}v\textup{d}x\textup{d}y$. This inner product does not depend on the choice of $U$, so in what follows, we drop the subscript.
We now specify to our case by taking $\Lambda=L=\mathbb{Z}^{n}$ and $\sigma(x,y)=x^{t}Sy$ for all $x,y\in V$. We also write $J_{S}$ for the integral Jacobi group in this case, i.e. $J_{S}:=\textup{SL}_{2}(\mathbb{Z})\rtimes(\mathbb{Z}^{n}\times\mathbb{Z}^{n})$.
@@
$\Gamma_{S,J}:=\left\{\begin{pmatrix}*&*\\ 0&D\end{pmatrix}\in\Gamma_{S}\mid D\in\textup{Mat}_{2}(\mathbb{Z})\right\}.$
@@
with $D^{*}:=D\left[\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}\right$] (cf. [eisenstein_thesis, page 44]). We denote by $\Gamma_{S,J}^{\bullet}:=\iota(J_{S})$. By [ajouz, Proposition 2.2.7], we have that the action of $J_{S}$ on $\mathbb{H}\times\mathbb{C}^{n}$ is given by
@@
where $j^{\prime}$ denotes the usual factor of automorphy for the action of $\textup{SL}_{2}(\mathbb{Z})$ on $\mathbb{H}$.
@@
is the same as the action of $J_{S}$ on $\mathbb{H}\times\mathbb{C}^{n}$. We now have the following Proposition regarding the fundamental domain of the action of $\Gamma_{S,J}^{\bullet}$ on $\mathcal{H}_{S}$, which will be useful later.
@@
$\mathcal{F}^{J}:=\left\{Z=(\omega,z,\tau)\in\mathcal{H}_{S}\mid(z,\tau)\in\mathcal{F},y_{1}y_{2}-\frac{1}{2}S[v]>0,\ -\frac{1}{2}\leq x_{1}\leq\frac{1}{2}\right\},$
where $\mathcal{F}$ is a fundamental domain for the action of $J_{S}$ on $\mathbb{H}\times\mathbb{C}^{n}$.
@@
Let $Z=(\omega,z,\tau)\in\mathcal{H}_{S}$. We can then pick $g\in\Gamma_{S,J}^{\bullet}$ such that $\iota^{-1}(g)\in J_{S}$ and that $Z^{\prime}:=g\langle Z\rangle=(\omega^{\prime},z,\tau)$ with $(z,\tau)\in\mathcal{F}$ and $\omega^{\prime}\in\mathbb{H}$ arbitrary. This follows from the fact that the actions are the same, as we have shown above. If now
@@
for $\lambda\in\mathbb{Z}^{n+2}$, then $T_{\lambda}\in\Gamma_{S}$ (cf. [eisenstein_thesis, Theorem 1.4.4]). We then have $T_{\lambda}\langle Z^{\prime}\rangle=Z^{\prime}+\lambda$ and so we can act with a suitable $T_{\lambda}$, so that $-1/2\leq x_{1}\leq 1/2$. Finally, the condition $y_{1}y_{2}-S[v]/2>0$ follows from the fact that $Z\in\mathcal{H}_{S}$ and the definition of $\mathcal{P}_{S}$ in Section 2. ∎
4. Dirichlet Series and Integral Representation
In this Section, we define the Dirichlet series of interest and give its integral representation in terms of an Eisenstein series of Klingen type.
@@ 4. Dirichlet Series and Integral Representation
$\mathcal{D}_{F,G}(s):=\sum_{m=1}^{\infty}\langle\phi_{m},\psi_{m}\rangle m^{-s}.$
@@
$\mathcal{D}_{F,G}(s)$ converges absolutely for $\textup{Re}(s)>k+1$ and represents a holomorphic function on this domain.
@@
with the constant depending only on $F,G$. Indeed, fix $(\tau,z)\in\mathbb{H}\times\mathbb{C}^{n}$ and write $\tau=u+iv,\textup{ }z=x+iy$. If $F(q)=\sum_{N=1}^{\infty}\phi_{N}(\tau,z)q^{N}$, with $q=e^{2\pi i\omega}$, we have by Cauchy’s integral formula that
$\phi_{N}(\tau,z)=\oint_{|q|=r}\frac{F(q)}{q^{N+1}}\textup{d}q,$
@@
$\phi_{N}(\tau,z)=\int_{0}^{1}F(Z)e^{-2\pi iN\omega}\textup{d}u^{\prime},$
for any $v^{\prime}>Q[y]/v$. But now $\left|F(Z)\right|\left(Q_{0}[\textup{Im}Z]\right)^{k/2}$ is bounded on $\mathcal{H}_{S}$ from [hauffe21, II, Lemma 3.28], say by a constant $C>0$. Therefore, after choosing $v^{\prime}=Q[y]/v+1/N$, we have
$\left|\phi_{N}(\tau,z)\right|\leq Ce^{2\pi}\int_{0}^{1}\left(Q_{0}[\textup{Im}Z]\right)^{-k/2}e^{2\pi NQ[y]/v}\textup{d}u^{\prime}=Ce^{2\pi}\left(\frac{v}{N}\right)^{-k/2}e^{2\pi NQ[y]/v}.$
@@
$E(Z,s):=\sum_{\gamma\in\Gamma_{S,J}\backslash\Gamma_{S}}\left(\frac{Q_{0}[\textup{Im}(\gamma Z)]}{\textup{Im}((\gamma Z)_{2})}\right)^{s},$
where for $Z=(\omega,z,\tau)\in\mathcal{H}_{S}$, we write $Z_{2}:=\tau$. Also, recall $Q_{0}=S_{0}/2$.
@@
$\gamma=\begin{pmatrix}*&*\\ 0&D\end{pmatrix},\textup{ }D\in\textup{SL}_{2}(\mathbb{Z}),$
(cf. [krieg_integral_orthogonal, Proposition 3]). Now, for $Z=(\omega,z,\tau)\in\mathcal{H}_{S}$, we have $\gamma\langle Z\rangle_{2}=D\langle\tau\rangle$, where the action on the right denotes the usual action of $\textup{SL}_{2}(\mathbb{Z})$ on $\mathbb{H}$ (cf. [eisenstein_thesis, page 115]). By [Borcherds_Products, Lemma 3.20], we have
@@
where again $j^{\prime}$ denotes the usual factor of automorphy for the action of $\textup{SL}_{2}(\mathbb{Z})$ on $\mathbb{H}$. But we have $j(\gamma,Z)=j^{\prime}(D,\tau)$ and hence $E(Z,s)$ is well-defined.
@@
But by the proof of [eisenstein_thesis, Theorem 3.1.1], the sum
@@
Let $F,G\in S_{k}(\Gamma_{S})$. For $Z\in\mathcal{H}_{S}$ and $s\in\mathbb{C}$ with $\textup{Re}(s)>n+2$, we have
$\langle F(Z)E(Z,s),G(Z)\rangle=\frac{1}{\#\textup{SO}(S;\mathbb{Z})}(4\pi)^{-(s+k-n-1)}\Gamma(s+k-n-1)\mathcal{D}_{F,G}(s+k-n-1),$
@@
$\textup{SO}(S;\mathbb{Z}):=\{g\in\textup{SL}_{n}(\mathbb{Z})\mid g^{t}Sg=S\},$
@@
Let $I(s):=\langle F(Z)E(Z,s)\textup{, }G(Z)\rangle$. If we denote by $\mathcal{Q}$ a fundamental domain for the action of $\Gamma_{S}$ on $\mathcal{H_{S}}$, by using the standard unfolding argument, we have for $\textup{Re}(s)>n+1$
$\displaystyle I(s)$ $\displaystyle=\int_{\mathcal{Q}}F(Z)\overline{G(Z)}\sum_{\gamma\in\Gamma_{S,J}\setminus\Gamma_{S}}\left(\frac{Q_{0}[\textup{Im}(\gamma Z)]}{\textup{Im}((\gamma Z)_{2})}\right)^{s}\left(Q_{0}[\textup{Im}Z]\right)^{k}\textup{d}^{*}Z$
$\displaystyle=\frac{1}{[\Gamma_{S,J}:\Gamma_{S,J}^{\bullet}]}\int_{\mathcal{F}^{J}}F(Z)\overline{G(Z)}\left(\frac{Q_{0}[Y]}{\textup{Im}(\tau)}\right)^{s}\left(Q_{0}[Y]\right)^{k-n-2}\textup{d}X\textup{d}Y,$
where $\mathcal{F}^{J}$ is a fundamental domain for the action of $\Gamma_{S,J}^{\bullet}=\iota(J_{S})$ on $\mathcal{H}_{S}$ and $Z=(\omega,z,\tau)=X+iY$. We note here
$\Gamma_{S,J}:\Gamma_{S,J}^{\bullet}]=\#\textup{SO}(S;\mathbb{Z})<\infty,$
because $S$ is positive definite (cf. [eisenstein_thesis, Section 1.7]). Hence, from Proposition 3.8, with the same notation as there, we get
$I(s)=\frac{1}{\#\textup{SO}(S;\mathbb{Z})}\int_{\mathcal{F}}\int_{y_{1}y_{2}-\frac{1}{2}S[v]>0}\int_{-1/2\leq x_{1}\leq 1/2}F(Z)\overline{G(Z)}y_{2}^{-s}\left(y_{1}y_{2}-\frac{1}{2}S[v]\right)^{k-n-2+s}\textup{d}X\textup{d}Y,$
where $\mathcal{F}$ is a fundamental domain of the action of $J_{S}$ on $\mathbb{H}\times\mathbb{C}^{n}$. We now write
${F(Z)=\sum_{m=1}^{\infty}\phi_{m}(\tau,z)e^{2\pi im\omega}},\,\,\,\displaystyle{G(Z)=\sum_{m=1}^{\infty}\psi_{m}(\tau,z)e^{2\pi im\omega}}.$
Using the fact that for integers $m_{1},m_{2}$, we have
$\int_{-1/2}^{1/2}e^{2\pi i(m_{1}-m_{2})x_{1}}\textup{d}x_{1}=\begin{cases}1&\textup{ if }m_{1}=m_{2}\\ 0&\textup{ otherwise}\end{cases},$
@@
$I(s)=\frac{1}{\#\textup{SO}(S;\mathbb{Z})}\int_{\mathcal{F}}\int_{y_{1}y_{2}-\frac{1}{2}S[v]>0}\sum_{m=1}^{\infty}\phi_{m}(\tau,z)\overline{\psi_{m}(\tau,z)}e^{-4\pi my_{1}}y_{2}^{-s}\times\\ \times\left(y_{1}y_{2}-\frac{1}{2}S[v]\right)^{k-n-2+s}\textup{d}X\textup{d}Y.$
@@
$I(s)=\frac{1}{\#\textup{SO}(S;\mathbb{Z})}\int_{\mathcal{F}}\int_{t=0}^{\infty}\sum_{m=1}^{\infty}\phi_{m}(\tau,z)\overline{\psi_{m}(\tau,z)}e^{-4\pi mt}e^{-2\pi m\frac{S[v]}{y_{2}}}y_{2}^{-s}\left(ty_{2}\right)^{k-n-2+s}\textup{d}t\textup{d}u\textup{d}v\textup{d}x_{2}\textup{d}y_{2}.$
@@
$\int_{t=0}^{\infty}e^{-4\pi mt}t^{k-n-2+s}\textup{d}t=\Gamma(s+k-n-1)(4\pi)^{-(s+k-n-1)}m^{-(s+k-n-1)}.$
Moreover, in this case, the inner product of Definition 3.6 reads as:
$\langle\phi_{m},\psi_{m}\rangle=\int_{\mathcal{F}}\phi_{m}(\tau,z)\overline{\psi_{m}(\tau,z)}y_{2}^{k-n-2}e^{-\frac{2\pi m}{y_{2}}S[v]}\textup{d}u\textup{d}v\textup{d}x_{2}\textup{d}y_{2}.$
@@
$I(s)=\frac{1}{\#\textup{SO}(S;\mathbb{Z})}(4\pi)^{-(s+k-n-1)}\Gamma(s+k-n-1)\mathcal{D}_{F,G}(s+k-n-1),$
@@
$\langle F(Z)E(Z,s),G(Z)\rangle=\frac{1}{\#\textup{SO}(S;\mathbb{Z})}(4\pi)^{-(s+k-n-1)}\Gamma(s+k-n-1)\mathcal{D}_{F,G}(s+k-n-1),$